a charge of 2.00 uC flows into the plates of a capacitor when it is connected to a 12.0 V battery. How much work was done in charging this capacitor?
Answer #1
W=V×q =12(2×10^-6) =2.4×10^-5 =24 ×10^-6
source: Eric Mazur phys1033
answered by: Mukhethwa
Answer #2
W=0.5V×q =0.5(12)(2×10^-6) =1.2×10^5 =12 uJ
source: Eric Mazur phys1033
answered by: Mukhethwa